• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name G.M. Urciuoli

    Log entry time 16:06:29 on June10,2005

    Entry number 145765

    keyword=Day shift report


    8:00 Shift starts

    Shit crew:

    G.M. Urciuoli SL
    R. Lindgren
    G. Chang

    Thick carbon target; Beam Current 50 uA

    Time Run Number N. of Events
    8:06 4476 1M
    8:41 4477 1M

    run 4478 junk


    9:30 Some problem with ROC4 that had to be reset

    9:45 MCC turned off the beam for an access.

    10:56 beam on again Rich has been turned on. Beam current 50uA

    10:56 run 4479 5.000 events Rich pedestal run.
    11:03 run 4480 2.000 events
    11:17 run 4481 200K events

    Beam Current 75 uA Rich still ON
    prescalers 375, 56, 68, 76, 1, 260, 1

    Time Run Number N. of Events
    11:25 4482 280k
    11:32 4483 1M

    11:50 The rich apparantely has some HV problem. An access is requested to check.

    12:31 run 4484: Cosmic run. 3k events
    Prescalers 65535, 65535, 1, 65535, 65535, 65535
    12:38 run 4485 Cosmic run 18k Events
    1337 run 4486 Cosmic run 19k Events

    Beam on again 50 uA prescalers 375, 56, 68, 76, 1, 260, 1
    Time Run Number N. of Events
    14:17 4487 80k

    14:25 Beam off again
    15:30 rigt Q1 has an Eath Leakege failure. We request an access.